# CubicML: Automated ML for Distributed ML Systems Co-design with ML Prediction of Performance

### v1

@@ Abstract
Scaling up deep learning models have been proven effective to improve intelligence of machine learning (ML) models, especially for industry recommendation models and large language models. The co-design of distributed ML systems and algorithms (to maximize training performance) plays a pivotal role for its success. As it scales, the number of co-design hyper-parameters grows rapidly which brings challenges to feasibly find the optimal setup for system performance maximization. In this paper, we propose CubicML which uses ML to automatically optimize training performance of distributed ML systems. In CubicML, we use a ML model as a proxy to predict the training performance for search efficiency and performance modeling flexibility. We proved that CubicML can effectively optimize training speed of in-house ads recommendation models and large language models at Meta.
@@ 3.2 ML Prediction of Distributed Training Performance of Large Language Models
We use training speed WPS (words/tokens per second) as the performance metric that CubicML targets to model. The search space includes configurations detailed in Appendix A.2.2, including Transformer model architecture configurations [21], distributed training co-design configurations [20] and system infra setup. Each configuration has a wide range of values. For example, the number of GPUs can be as small as $8$ and up to $16,384$; the sequence length ranges between $2048$ and $131,072$. The predictor training details are in Appendix A.1.2. We accumulated $568$ LLM jobs over time during LLM development. We use $145$ examples as validation dataset and the rest ($423$ examples) as training data. The predicted WPS versus actual WPS is plotted in Figure 3, where we performed different training-validation split to simulate three different real-world use cases:
  • using random split which simulates the use case where CubicML can sample and profile configurations in real-time (i.e. online) for performance optimization as the ML model and systems evolves;
@@ A.2.1 Ads Recommendation Models
  • Storage reservation policy: percentage of HBM reserved for dense paramemers and runtime memory, ranging from 0.77 to 0.85, each step is 0.01.
@@ A.2.2 Large Language Models
  • Distributed training co-design configurations [20], such as parallelism strategies (tensor parallelism, pipeline parallelism, context parallelism, data parallelism) and their configurations, precision formats (FP8 and BF16), etc
@@
Each configuration has a wide range of values. For example, the number of GPUs can be as small as $8$ and up to $16,384$; the sequence length ranges between $2048$ and $131,072$.

### v2

# CubicML: Automated ML for Large ML Systems Co-design with ML Prediction of Performance

@@ Abstract
Scaling up deep learning models has been proven effective to improve intelligence of machine learning (ML) models, especially for industry recommendation models and large language models. The co-design of large distributed ML systems and algorithms (to maximize training performance) plays a pivotal role for its success. As it scales, the number of co-design hyper-parameters grows rapidly which brings challenges to feasibly find the optimal setup for system performance maximization. In this paper, we propose CubicML which uses ML to automatically optimize training performance of large distributed ML systems. In CubicML, we use an ML model as a proxy to predict the training performance for search efficiency and performance modeling flexibility. We proved that CubicML can effectively optimize training speed of in-house ads recommendation models with $73$ billion parameters and large language models up to 405 billion parameters at Meta.
@@ 3.2 ML Prediction of Distributed Training Performance of Large Language Models
We use training speed WPS (words/tokens per second) as the performance metric that CubicML targets to model. The search space includes configurations detailed in Appendix A.2.2, including Transformer model architecture configurations [21], distributed training co-design configurations [20] and system infra setup. Each configuration has a wide range of values. For example, the number of GPUs can be as small as $8$ and up to $16,384$; the sequence length ranges between $2048$ and $131,072$; the number of parameters ranges from $27$ million up to $405$ billion. The predictor training details are in Appendix A.1.2. We accumulated $568$ LLM jobs over time during LLM development. We use $145$ examples as validation dataset and the rest ($423$ examples) as training data. The predicted WPS versus actual WPS is plotted in Figure 3, where we performed different training-validation split to simulate three different real-world use cases:
  • using random split which simulates the use case where CubicML samples and profiles configurations in real-time (i.e. online) for optimization as model and systems evolve;
@@ A.2.1 Ads Recommendation Models
  • Storage reservation policy:
    • FixedPercentage: A fixed percentage of HBM reserved for dense paramemers and runtime memory, ranging from 0.77 to 0.85, each step is 0.01.
    • MemoryBalanced: The embedding sharding planner will find the minimal HBM used for emebdding tables, leave rest for dense parameters and runtime memory.
@@ A.2.2 Large Language Models
  • Distributed training co-design configurations [20], such as which parallelism strategies (tensor parallelism, pipeline parallelism, context parallelism, data parallelism) should be selected and their configurations (e.g. group sizes), precision formats (FP8 and BF16), etc
@@
Each configuration has a wide range of values. For example, the number of GPUs can be as small as $8$ and up to $16,384$; the sequence length ranges between $2048$ and $131,072$; the number of parameters ranges from $27$ million up to $405$ billion.